Calculate the amount of Calcium chloride dihydrate in grams needed to make a 250 mL solution with a final Normality of 2,000 meq/L. Round answer to the nearest hundredth. 


1
Expert's answer
2021-03-22T06:31:51-0400

Cn = n/V

n = m/e

e = M/valency factor

M (Ca(OH)2) = 74.1 g/mol

e (Ca(OH)2) = 74.1/2 = 37.05

2,000 meq/L = 2 eq/L

n (Ca(OH)2) = Cn x V = 2 x 0.25 = 0.5 n

m (Ca(OH)2) = n x e = 0.5 x 37.05 = 18.5 g